{5x + 3y= 7
{x - 6y = 8

Answer by rfer(16322)

Graph the following:
X-6Y=8 5X+3Y=7
(0,-1 1/2) (2,-1)
(2,-1) (1,2/3)
(8,0) (4,-4 1/3)
The common point or intersection is (2,-1)